In Okunland, a country whose economy operates according to Okun's law, real GDP equals $7,520 billion, potential GDP equals $8,000 billion, and the actual unemployment rate is 8 percent. What is the natural rate of unemployment in Okunland? A) 5 percent B) 6 percent C) 11 percent D) 2 percent
Answer #1

Gap = 8000-7520 = 480

As real GDP is less than potential GDP so there exist a recessionary gap

output gap = 480/8000*100 = 6%

As per Okun law, each percentage point of unemployment above its natural rate is equal to 2% point of the recessionary gap so,

As the output is 6% below its potential output so the actual unemployment is 3% above its natural rate of unemployment

Therefore natural rate of unemployment is 8-3 = 5%

option(A)